OpenStudy (anonymous):

The length of a rectangle is increasing at a rate of 10 cm/s and its width is increasing at a rate of 6 cm/s. When the length is 12 cm and the width is 10 cm, how fast is the area of the rectangle increasing? Found the answer. It is 172square cm/sec now need to find How fast is the perimeter of the rectangle increasing?

OpenStudy (raden):

let the length be l, width w l = 10 ----> l' = 12 w = 6 ---> w' = 10 if A = lw then, A' = l'*w + l*w' A' = 12 * 6 + 10 * 10 = 72 + 100 = 172

OpenStudy (anonymous):

Yeah i found 172 need to find how fast is the PERIMETER of the rectangle increasing